OpenStudy (anonymous):

5-v 3v-8 ---- - ------ = v-7 7-v

OpenStudy (anonymous):

i got 2v-3 ---- v-7

OpenStudy (anonymous):

since the denominators are the same we can just subtract the numerators to get (5-v)-(3v-8)=5-v-3v+8 =-4v+13

OpenStudy (anonymous):

thus your final answer is (-4v+13)/(v-7)

OpenStudy (anonymous):

\[\frac{5-v}{v-7}+\frac{3v-8}{v-7}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

v-7

OpenStudy (anonymous):

7-v